theyadoreshayyy:

Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3,136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percentage. Provide an explanation and proof for your answer to receive full credit.

xxemilianaxx:

\[v = \frac{ 1 }{ 3 } BH\]

theyadoreshayyy:

i need the work for it though

xxemilianaxx:

I know that's why were going step by step

theyadoreshayyy:

okay

xxemilianaxx:

So since b x h is the volume of the entire rectangular prism with the height h \[v = \frac{ 1 }{ 3 } BH\] Can you input the rest?

xxemilianaxx:

You need to find out how much larger the second one is to this smaller one. So, to do it into a percent, Divide the smaller one by the larger one and then multiply it by 100 to get it in percentage. \[28\div 3136\] Doing so will make it a decimal so if you multiply it by 100 (as you know) that makes it a percentage

theyadoreshayyy:

i got 0.89285714285

xxemilianaxx:

\[0.89285714285\] Should be what you get Volume of pyramid B is \[ 8\] times or \[12.5% \] bigger than pyramid A.

xxemilianaxx:

Tis not really exactly 1pecent of the volume so, It make's more sense because 28 is smaller than 3136

supie:

Asked & Answered here: https://questioncove.com/study#/updates/53c9961ae4b0bff525b153c3

surjithayer:

volume of prism A\[=\frac{ 1 }{ 3 }\times base~ area \times height=\frac{ 1 }{ 3 }\times 14\times 14\times 6=392 ~\ inch^3\] volume of pyramid B=3,136 cubic inch \[\frac{ volume~of~pyramid~B }{ volume~of~pyramid~A }=\frac{ 3136 }{ 392}=8\] so pyramid B is8 times larger in volume than volume of pyramid A